OpenStudy (anonymous):

15a^b/-7xy^2 divided by -3ab/14xy

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

\[\frac{15a^b}{-7xy^2} *\frac{14xy}{-3ab}\] \[=\frac{10a^b}{yb}\]

OpenStudy (anonymous):

Looking at this, shouldn't the answer be: \[(10z ^{b-1})/yb\] ? I see your bottom a go away without the top one being impacted.

OpenStudy (anonymous):

That should be 10a^(b-1)

OpenStudy (anonymous):

Thank you for the correction

OpenStudy (anonymous):

I'm rusty on some of the algebra, but I think the rule is: \[x ^{m}/x ^{n}=x ^{m-n}\]